# A new machine learning framework for occupational accidents forecasting with safety inspections integration

### v1

@@ 7 Discussion
The proposed framework aims at supporting proactive accident prevention by providing weekly assessments of risk levels based on recurrent safety inspections. After analyzing feedbacks from safety inspections, the system estimates the probability of an accident for each day. These probabilities are then compared to a calibrated threshold: if the probability exceeds this threshold, the day is flagged as at risk. A week is classified as risky if at least one day exceeds the threshold.
@@
This setup enables safety teams to anticipate high-risk periods. In practice, the model can be used at the beginning of each week to generate a brief report indicating whether the upcoming week is considered safe or risky, and which specific days require attention. For example, if Week 4 is flagged as risky due to to day 1 and 2, the safety officer may schedule targeted actions such as site visits, safety briefings, or specific inspections on those days. Thanks to its time-based design, the model can be regularly updated with newly reported field data. It can also be easily integrated into a prevention dashboard with automatic updates (e.g., every Monday) to support weekly planning. The more frequently and consistently field reports are submitted, the more robust the model becomes capturing subtle shifts in weak signals and risky behaviors over time. Despite these advantages, the approach also presents certain limitations. Aggregating data at the weekly level may sometimes under represent risks spread across several moderately risky days. For instance, three consecutive days with medium-level probabilities might not trigger any alert, whereas a single day with a high probability could result in the whole week being labeled as risky. This simplification should be kept in mind when interpreting results. Moreover, the poor performance of some traditional models may stem from poorly calibrated probability outputs. If the predicted scores do not accurately reflect reality, it becomes difficult to set effective thresholds to differentiate safe from risky periods. Improving probability calibration enhance the reliability of the alerts. Another limitation lies in the forecasting strategy used, known as DirRec, which combines direct and recursive predictions. While this method helps capture temporal dependencies, it can also amplify errors: an inaccurate prediction early in the week may propagate to later days, degrading the overall weekly forecast quality. Several directions can be explored to improve the system. Better-calibrated probabilistic models could be developed, decision thresholds could be tailored to specific departments, and more advanced architectures, such as sequence-to-sequence models, could be considered. Additionally, the textual content of hazardous situation reports, currently summarized through quantitative indicators such as severity or frequency, could be leveraged more thoroughly.Moreover, Semantic analysis of descriptions (e.g., using natural language processing techniques) could extract richer and potentially predictive signals, while capturing the specific context of the reported events.
@@ 8 Conclusion
This study introduced a generic and operational framework for short-term forecasting of occupational accidents based on binary time series. Using proactive data from safety inspections, the model dynamically predicts daily accident probabilities and provides a weekly classification of risk. The proposed approach demonstrated its effectiveness in identifying both risky and safe periods, particularly through the use of an LSTM model, which outperformed classical machine learning methods in evaluated series.
Due to its temporal structure and weekly aggregation, this framework is well suited to support real-time prevention strategies in industrial settings. It can be integrated into a safety dashboard and updated regularly to help decision makers plan targeted actions. The ability to anticipate high-risk weeks opens new avenues for more proactive, data-driven safety management.
Future improvements could focus on better probability calibration, more robust modeling architectures, and the incorporation of semantic information from textual safety reports. More broadly, this work shows the potential of time series forecasting and machine learning to transform occupational safety into a more predictive and preventive discipline.
Acknowledgements
The authors would like to acknowledge the financial support provided by the French government through a grant managed by the Agence Nationale de la Recherche (ANR) as part of the France 2030 investment plan (grant number ANR-23-EXMA-0001).
CRediT authorship contribution statement
Aho Yapi: Writing – review & editing, Writing – original draft, Software, Methodology, Formal analysis, Conceptualization. Pierre Latouche: Writing – review & editing, Methodology, Supervision. Arnaud Guillin: Writing – review & editing, Methodology, Supervision. Yan Bailly: Writing – review & editing, Methodology, Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could appear to influence the work reported in this paper.
@@ Appendix B Visualization of binary time series and rate evolution graph of ITW et ExW series
Appendix C Weekly accident risk prediction and comparison with observed accident counts on the ITW and ExW series with LSTM model
Figure 8: Weekly accident risk prediction and comparison with observed accident counts on the ITW series with LSTM model.
Figure 9: Weekly accident risk prediction and comparison with observed accident counts on the ExW series with LSTM model.

### v3

@@ 8 Discussion
The proposed framework aims at supporting proactive accident prevention by providing weekly assessments of risk levels based on recurrent safety inspections. After analyzing feedbacks from safety inspections, the system estimates the probability of an accident for each day. These probabilities are then compared to a calibrated threshold: if the probability exceeds this threshold, the day is flagged as at risk. A week is classified as risky if at least one day exceeds the threshold. The framework is model-agnostic: preprocessing, feature engineering, validation protocol and scoring procedures are applied identically across model families, enabling fair comparisons and easy substitution of prediction models. In practical terms, the output is designed as decision support rather than automatic enforcement: the threshold can be tuned to favor either recall (capture more risky weeks with more false alarms) or precision (fewer alerts with higher confidence), reflecting each site’s prevention policy and resource constraints. Alerts translate into simple playbooks for prevention teams (e.g., schedule a targeted walk-down or conduct a focused toolbox talk on the flagged day). This makes the signal immediately actionable for weekly planning. Rather than aggregating all accidents into a single binary label, we concentrate on one specific accident category defined in Appendix A and apply the same weekly pipeline to this category only. The model produces category-specific daily probabilities; a day is flagged as at risk for that category when its probability exceeds the calibrated threshold, and a week is deemed risky for that category if at least one day is flagged. This category focus yields status labels that directly reflect the expected occurrence of that accident type and enables more targeted prevention actions. The weekly roll-up integrates naturally with leading indicators already tracked by HSE (inspection counts, hazardous situations, improvement actions): combining a risky week flag with weak-signal trends helps prioritize scarce resources and improves the timeliness of preventive actions.
@@
This setup enables safety teams to anticipate high-risk periods. In practice, the model can be used at the beginning of each week to generate a brief report indicating whether the upcoming week is considered safe or risky, and which specific days require attention. For example, if Week 4 is flagged as risky due to to day 1 and 2, the safety officer may schedule targeted actions such as site visits, safety briefings, or specific inspections on those days. Thanks to its time-based design, the model can be regularly updated with newly reported field data. It can also be easily integrated into a prevention dashboard with automatic updates (e.g., every Monday) to support weekly planning. The more frequently and consistently field reports are submitted, the more robust the model becomes capturing subtle shifts in weak signals and risky behaviors over time. Despite these advantages, the approach also presents certain limitations. Aggregating data at the weekly level may sometimes under represent risks spread across several moderately risky days. For instance, three consecutive days with medium-level probabilities might not trigger any alert, whereas a single day with a high probability could result in the whole week being labeled as risky. This simplification should be kept in mind when interpreting results. Moreover, the poor performance of some traditional models may stem from poorly calibrated probability outputs. If the predicted scores do not accurately reflect reality, it becomes difficult to set effective thresholds to differentiate safe from risky periods. Improving probability calibration enhance the reliability of the alerts. Several directions can be explored to improve the system. Better-calibrated probabilistic models could be developed; decision thresholds could be tailored to specific departments; more advanced architectures, such as sequence-to-sequence models, could be considered; and global forecasting could help capture cross-site dependencies. Beyond accuracy, future work will report cost-sensitive analyses contrasting the expected cost of a missed risky week versus the operational cost of investigating a false alarm, which is central to prevention impact. Additionally, the textual content of hazardous-situation reports, currently summarized through quantitative indicators such as severity or frequency, could be leveraged more thoroughly. Semantic analysis of descriptions (e.g., using natural language processing techniques) could extract richer and potentially predictive signals while better capturing the specific context of reported events.
@@ 9 Conclusion
This study introduced a model-agnostic and operational framework for short-term forecasting of occupational accidents based on binary time series. Using proactive data from safety inspections, the model dynamically predicts daily accident probabilities and provides a weekly classification of risk. The proposed approach demonstrated its effectiveness in identifying both risky and safe periods, particularly through the use of an LSTM model, which outperformed classical machine learning methods in evaluated series. From a prevention perspective, the weekly flag and day-level probabilities offer a simple, auditable signal that can be embedded in standard routines (weekly safety meetings, planning boards, shift handovers) to prioritize inspections and briefings precisely when they are most needed. Because the pipeline is model-agnostic and thresholds are tunable, sites can adopt operating modes that reflect their risk appetite and seasonality, while keeping a consistent validation protocol across algorithms.
Due to its temporal structure and weekly aggregation, this framework is well suited to support real-time prevention strategies in industrial settings. It can be integrated into a safety dashboard and updated regularly to help decision makers plan targeted actions. The ability to anticipate high-risk weeks opens new avenues for more proactive, data-driven safety management. Future deployments will include calibration monitoring and drift checks, horizon-specific summaries for short maintenance windows, and category-aware dashboards that link each risky week to concrete action checklists. We anticipate these enhancements will further translate predictive accuracy into measurable reductions in incident rates and improved timeliness of preventive actions.
@@ Appendix B Visualization of binary time series and rate evolution graph of ITW et ExW series
Appendix C Weekly accident risk prediction and comparison with observed accident counts on the ITW and ExW series
Figure 10: Weekly accident risk prediction and comparison with observed accident counts on the ITW series with LSTM-MIMO model.
Figure 11: Weekly accident risk prediction and comparison with observed accident counts on the ExW series with decision-tree model.
Appendix D Hyperparameters grids and selected settings
Table 9 summarizes the grids and the settings selected by grid search for a weekly horizon H=7H{=}7. The lag depths ($d_{y}$ and $d_{c}$) are tuned over $\{7,14,28\}$. Implementations rely on scikit-learn and native libraries. For each dataset (ITW, ITW-d1, EXW), the table reports the best hyperparameters together with the selected lags $(d_{y},d_{c})$ and the decision threshold $\tau$.
Classifier Hyperparameters grid ITW ITW-d1 EXW
XGBoost learning_rate = [0.01, 0.05, 0.1] max_depth = {3, 5, 10} n_estimators = [50, 100, 200] reg_alpha = [0, 0.5, 1] reg_lambda = [1, 3, 10] scale_pos_weight = [1, 3.7, 5, 10, 20.88] lr=0.1, n=100, depth=5, ra=0.5, rl=1.0, spw=3.7; H=7, dy=14, dc=14, $\tau$=0.5 lr=0.1, n=100, depth=3, ra=0.5, rl=3.0, spw=20.88; H=7, dy=28, dc=28, $\tau$=0.4 lr=0.1, n=50, depth=3, ra=0.0, rl=3.0, spw=1; H=7, dy=14, dc=28, $\tau$=0.2
Logistic regression C = [0.01, 0.1, 1, 5, 10] penalty = {l1, l2} class_weight = {None, balanced} C=0.1, pen=l1, cw=None; H=7, dy=28, dc=14, $\tau$=0.5 C=0.1, pen=l2, cw=None; H=7, dy=14, dc=14, $\tau$=0.25 C=5, pen=l1, cw=None; H=7, dy=28, dc=28, $\tau$=0.1
LightGBM learning_rate = [0.01, 0.05, 0.1] n_estimators = [50, 100, 200] reg_alpha = [0, 0.5, 1] reg_lambda = [1, 3, 10] scale_pos_weight = [1,3.7,10, 20.88] lr=0.05, n=50, ra=0.5, rl=3.0, spw=3.7; H=7, dy=14, dc=14, $\tau$=0.45 lr=0.05, n=50, ra=0.5, rl=3.0, spw=20.88; H=7, dy=14, dc=28, $\tau$=0.2 lr=0.05, n=100, ra=0.5, rl=1.0, spw=10; H=7, dy=14, dc=14, $\tau$=0.35
Decision tree criterion = {gini, entropy} max_depth = {5, 10, 15, None} min_samples_split = [2, 5, 10] min_samples_leaf = [1, 3, 5] class_weight = {None, balanced} crit=entropy, depth=None, split=2, leaf=5, cw=None; H=7, dy=28, dc=14, $\tau$=0.6 crit=entropy, depth=15, split=2, leaf=5, cw=balanced; H=7, dy=14, dc=14, $\tau$=0.8 crit=entropy, depth=10, split=2, leaf=5, cw=balanced; H=7, dy=14, dc=14, $\tau$=0.75
MLP hid_layer_sizes = {4, 16 32,64,128,64} alpha = [1e-4, 1e-3, 1e-2] batch_size = {32, 64} max_iter = {100, 200} hls=(4,), alpha=1e-3, bs=32, iters=200; H=7, dy=14, dc=28, $\tau$=0.75 hls=(4,), alpha=1e-3, bs=32, iters=100; H=7, dy=28, dc=28, $\tau$=0.3 hls=(128,64), alpha=1e-3, bs=64, iters=200; H=7, dy=14, dc=14, $\tau$=0.25
HistGradientBoosting learning_rate = [0.01, 0.05, 0.1] max_depth = {3, 5, 6, 10, None} max_iter = {50, 100, 200} l2_regularization = [0.0, 0.01, 0.1] n_iters = [10,50,100] lr=0.1, depth=6, iters=50, l2=0.01; H=7, dy=28, dc=14, $\tau$=0.25 lr=0.1, depth=6, iters=100, l2=0.01; H=7, dy=28, dc=14, $\tau$=0.15 lr=0.1, depth=None, iters=100, l2=0.01; H=7, dy=14, dc=28, $\tau$=0.3
Random forest n_estimators = [50, 100, 200] max_depth = {5, 10, 20, None} min_samples_split = [2, 5, 10] min_samples_leaf = [1, 3, 5] class_weight = {None, balanced} n=50, depth=None, split=2, leaf=3, cw=balanced; H=7, dy=14, dc=14, $\tau$=0.4 n=50, depth=None, split=5, leaf=3, cw=balanced; H=7, dy=14, dc=28, $\tau$=0.15 n=50, depth=5, split=2, leaf=5, cw=balanced; H=7, dy=14, dc=14, $\tau$=0.45
Table 9: Best hyperparameters identified by grid search for machine learning models.
Table 10 lists the grids and selected settings for sequence models at H=7H{=}7. Architectures are implemented in PyTorch via Darts. Since binary heads are not provided natively, we added a sigmoid output layer and train with a binary cross-entropy loss so that models output calibrated class-1 probabilities. The sequence length seq_len is tuned over $\{14,21,28\}$. For each dataset, the table reports the chosen architecture settings along with the batch size (bs), number of epochs (ep), and the selected threshold $\tau$.
Classifier Hyperparameters grid ITW ITW-d1 EXW
LSTM–MIMO seq_len = [14, 21, 28] hidden_dim = [64, 128] n_rnn_layers = [1, 2] dropout = [0.0, 0.2, 0.5] activation = {relu,tanh} seq=14; H=7; act=relu, hid=128, layers=1, dr=0.5; bs=64, ep=30; $\tau$=0.2 seq=14; H=7; act=relu, hid=128, layers=2, dr=0.0; bs=128, ep=30; $\tau$=0.1 seq=28; H=7; act=relu, hid=64, layers=2, dr=0.5; bs=128, ep=75; $\tau$=0.1
LSTM–Seq2Seq seq_len = [7, 14,28] hidden_dim = [64, 128] n_layers = [1, 2, 3] dropout = [0.0, 0.2, 0.5] activation = {relu,tanh} seq=14; H=7; hid=64,act=relu, layers=2, dr=0.5; bs=64, ep=50; $\tau$=0.95 seq=28; H=7; hid=64,act=relu, layers=1, dr=0.5; bs=64, ep=30; $\tau$=0.2 seq=14; H=7; hid=128, layers=2, dr=0.5; bs=32, ep=30; $\tau$=0.85
TCN seq_len = [7, 14,28] num_filters = [16, 32, 64] kernel_size = [3, 5] dilation_base = [2, 3] dropout = [0.0, 0.2, 0.5] weight_norm = {True, False} seq=14; filt=64, k=3, dil=3, dr=0.5, wn=False; bs=64, ep=50; $\tau$=0.55 seq=28; filt=64, k=5, dil=3, dr=0.2, wn=False; bs=64, ep=50; $\tau$=0.65 seq=14; filt=64, k=5, dil=3, dr=0.5, wn=True; bs=32, ep=50; $\tau$=0.7
TFT seq_len = [7, 14,28] hidden_size = [32, 64] attn_heads = [2, 4, 8] dropout = [0.0, 0.2, 0.5] seq=14; hid=32, heads=4, dr=0.1; bs=64, ep=30; $\tau$=0.50 seq=14; hid=64, heads=4, dr=0.1; bs=64, ep=30; $\tau$=0.40 seq=28; hid=32, heads=4, dr=0.1; bs=64, ep=30; $\tau$=0.30
Table 10: Best hyperparameters identified by grid search for deep learning models.